OpenStudy (anonymous):

A baseball goes from zero to 39 m/s in 0.197 s. What is its average acceleration? Answer in units of m/s 2

OpenStudy (sidsiddhartha):

Average accleration = change in velocity / change in time =(39-0)/0.197=39/0.197

OpenStudy (sidsiddhartha):

so what's ur answer?@helpme305